# Distinguishing Calabi-Yau Topology using Machine Learning

### v1

@@ Abstract
Keywords: AI-assisted mathematics, Machine learning, Calabi-Yau manifolds, String compactification
@@ 1 Introduction
In the realm of contemporary scientific inquiry, the application of deep learning has emerged as a transformative force, endowing researchers with novel tools and methodologies to explore intricate scientific phenomena. While AI has been transformative in the experimental sciences for decades, its revolution in theoretical and mathematical sciences have been a relatively more recent emergence [30, 44, 15, 48]; see reviews and prospects in [31, 49, 32, 8, 34, 28, 33]. A programme to machine-learn various structures in mathematics by looking at pure mathematical data had been launched over the past seven years, and have blossomed to include algebraic geometry [30, 31], algebra and representation theory [35, 19], graph theory and combinatorics [39], knot theory [17, 29, 19], number theory [1, 36], symbolic computation [47, 21, 45], etc.
@@
Calabi-Yau manifolds are a protagonist in modern mathematics, residing at the interface between algebraic and differential geometry, mathematical physics (especially string theory), topology, number theory, and dynamical systems. These are Kähler manifolds that admit zero Ricci curvature. The simplest example is one that is familiar to the beginning student: the torus $T^{2}=S^{1}\times S^{1}$ can be complexified into a Riemann surface, which is in turn trivially Kähler and furthermore is Ricci-flat. The name “Calabi-Yau” originated in the Fields-winning proof by one of the authors (STY) [52] that settled the Calabi Conjecture [12], and was first coined by physicists [14].
@@
In searching for the Standard Model of particle physics within string theory compactifications, one of the first data-bases in modern geometry was created; this was the so-called CICYs, for “Complete Intersection Calabi-Yau threefolds” [13, 26, 27] (for pedagogical introductions, see the classic in [41], as well as a recent treatment [31]). These are Calabi-Yau manifolds of complex dimension three, realized as algebraic varieties embedded as complete intersections in products of projective spaces. From a data-scientific point of view, they are rather straight-forwardly represented as matrices of non-negative integers, which encode the multi-degree information of the homogeneous polynomials that define the variety. It was natural, therefore, that the first machine-learning experiments in pure mathematics was in consideration on this data-set [30]. Indeed, despite its rich properties, the precise nature of the Calabi-Yau manifold often proves challenging to intuitively grasp and compute, presenting an intriguing and demanding avenue for the application of deep learning.
@@
Now, one should bear in mind that in topology, characteristics that describe the structure of manifolds, such as the classic Betti numbers which describe the closed cycles, are often intricately and closely linked with the data representing these manifolds. This makes the precise derivation of these characteristics quite challenging. The analytic mathematical formulas for these properties are largely unexplored  [^1] Due to moduli-dependence of the such quantities as cohomology groups – unlike their alternating sums which give rise to Euler characteristics that can be more readily computed by index theorem – are very difficult to calculate. See [16] for a recent successful extraction of exact formulae for surfaces guided by machine-learning., and their computation relies on algorithms that are both complex and time-consuming. Therefore, the application of deep learning in the field of topology poses significant challenges. This paper is dedicated to utilizing the capabilities of deep neural networks to predict some key properties of the Calabi-Yau manifold, in order to explore the potential applications of deep learning within the realm of mathematics and theoretical science.
@@
Henceforth, we focus on CICY threefolds, which have been a pivotal class of manifolds in the context of string model building and, as mentioned, one of the first data-set in pure mathematics to enter the age of data [13] and the era of machine-learning [30]. Since then, there have been a host of activity on machine-learning of properties of CICYs. Neural network methodologies were applied to predict the Hodge numbers of CICY three-folds, favorability, and discrete symmetries [10, 11], indicating that they outperform classical statistical methods such as Support Vector Machines (SVM). Subsequently, the prediction accuracy of Hodge numbers were increased to over 99% by leveraging the Inception network model [22, 23], and more recent comparative study, with impressive accuracies, was performed in [43]. Other related directions have included CICY fourfolds [38, 25] and generalized CICYs [18] (see review in [24]), numerical metrics [7, 20, 3, 4, 6], weighted CICYs [9, 40, 46], distinguishing elliptic fibrations [37] and topologies [42], etc. This expansion reflects a continued effort to broaden the scope of neural network applications in predicting geometric properties, highlighting the versatility of these models in handling increasingly complex mathematical structures.
@@
Now, most of the work on machine-learning of CICY topological invariants have focused on Hodge numbers, whereas, as we will soon see, more refined invariants involve Chern classes and intersection numbers. This is an important issue because the “topological type” of a (smooth, compact, and simply connected) Calabi-Yau threefold, due to the extension of a theorem of Wall [51], consists of a pair of Hodge numbers $(h^{1,1},h^{2,1})$, as well as (when expressed in an appropriate basis of the Kähler classes) the second Chern class $(c_{2})_{r}$ and the triple-intersection numbers $d_{rst}$ of the curve classes. It is a standing conjecture of one of the authors (STY), that in any complex dimension $n$, the possible topological types of a Calabi-Yau $n$-fold is finite.
@@
In sum, we hope to explore and expand the potential of deep learning to predict more critical properties of the CICY three-folds. In this study, we seek to predict the triple intersection numbers of CICY three-folds, combining ideas from [30, 22, 42], by an inception network V3 model. By amalgamating the nonlinear feature extraction capabilities of deep learning with the geometric properties of manifolds, we aim to enhance the deep learning model’s understanding of the CICY three-folds.
@@ Acknowledgments
YHH is grateful to the STFC, UK for grant ST/J00037X/2 and a Leverhulme project grant. ZGY has been supported by Singapore Ministry of Education Tier 2 grant A-8001562-00-00 and Tier 1 grant A8000987-00-00 at the National University of Singapore. ZGY also thanks the London Mathematical Society and the Royal Statistical Society for their partial support of his research trip to the UK.
@@ 2.1 Dataset
CICY manifolds are realized as a complete intersection of polynomials in a product of complex projective spaces [13, 26]. What this means is the following. Consider the ambient space as the product $X=\mathbb{C}P^{n_{1}}\times\cdots\times\mathbb{C}P^{n_{m}}$, a CICY three-fold $M^{3}$ is embedded as $k$ polynomials, the complete intersection condition is that
@@
Thus, each CICY is represented by an $m\times k$ matrix of non-negative (and indeed largely sparse) integers.
Now, $M^{3}$ is a Calabi-Yau three-fold, so it has vanishing first Chern class. This conveniently translates to (using adjunction, see e.g., [41, 31]) the condition that row sums to one less than dimension of the corresponding projective factor:
@@
The CICY three-folds were classified in [13], using the then state-of-the-art computing. Up to trivial permutation equivalence and also simple equivalence due to birational transformations, there are 7890 configuration matrices, giving rise to perhaps the first “big” database in geometry. Today, this database was updated in [5] and maintained at
@@
descends from the ambient $X$, and and we have corresponding to $k$ cohomology classes $q_{1}^{1}x_{1}+\cdots+q_{m}^{1}x_{m},\ \cdots,\ q_{1}^{k}x_{1}+\cdots+q_{m}^{k}x_{m}$. It should be emphasized that this implies that while $h^{1,1}(M^{3})$ come at least from the $m$ projective factors, by no means they are equal in general, as extra Kähler classes arise from the restriction. In the case of equality $h^{1,1}(M^{3})=m$, when all divisors of $M^{3}$ descend from the simple ambient space $X$, we refer to $M^{3}$ as favourable. Of the original 7890 CICY threefold configuration matrices [13], there are 4896 favorable geometries and 2994 unfavorable geometries [2]. Furthermore, new and favourable descriptions of 2946 unfavorable CICYs were presented therein. The favorable list can be obtained from http://www1.phys.vt.edu/cicydata/.
@@
Theorem 2.1 (C. T. C. Wall [51]).
@@
[…]
@@
This list of integer (tensors) is the topological datum for out Calabi-Yau threefold.
@@ 2.2 CICY Intersections
where we have also recorded its relation to the Hodge numbers, as a consequence of the index theorem. Alos, by Poincaŕe duality, $H^{4}(M,\mathbb{Z})\cong H^{2}(M,\mathbb{Z})$, thus second Chern class $c_{2}$ can be labelled by $h^{1,1}$ as
$[c_{2}(M)]_{t}=\sum_{r,s}[c_{2}(M)]_{rs}d_{rst}.$ (10)
@@
In general the computation of $(d_{rst},c_{2})$ is applicable for favorable geometries, we will design the computer algorithm later. On the other hand, it is usually hard to determine whether two set of data $(d_{rst},c_{2})$ corresponding to the same topology, i.e., to determine whether the difference between them corresponding merely to a change of basis of $H^{2}(M,\mathbb{Z})$. Thus, we adopt a family of divisibilitiy invariants (see §8.1 of [41]) up to change of basis:
@@
will be used to distinguish CICY topology. Since Hodge number prediction had been the subject of the literature since 2017, we will here focus on $(d_{1},d_{2},d_{3},d_{p})$.
@@ 2.3 Computer Algorithm for $d_{rst}$
    where $\sigma=(\sigma_{1},\sigma_{2},\cdots,\sigma_{m})$ is the partition of set $[k]=\{1,\cdots,k\}$, $q_{i}^{\sigma_{i}}:=\prod_{j\in\sigma_{i}}q_{i}^{j}$. $|\sigma_{i}|=n_{i},\ i\neq r,s,t$; $|\sigma_{i}|=n_{i}-1$ otherwise.
@@
    Second, we construct a $k\times k$ square matrix $A_{rst}(i,j)$ by adding $n_{i}-1$ copies of the $i$-th row to $A$ if $i\neq r,s,t$. Otherwise, we add $n_{i}-2$ copies of the $i$-th row to $A$. Note that the subscript $rst$ is just to show that the matrix depends on the indices of the triple intersection numbers $d_{rst}$ and the row-column indices that actually define the matrix are denoted as $(i,j)$.
@@
1:function detm($A$)
2:     if number of rows in A $=1$ then
@@
4:     else if number of rows in A $=2$ then
@@ 3.1 Inception Network model
To construct a robust architecture, we leverage the Inception Neural Network Model from GoogleNet [50]. This deep convolutional neural network architecture employs inception modules, a unique design that captures features across various spatial scales and complexities concurrently, departing from the sequential approach of traditional Convolutional Neural Networks. The distinctive feature of Inception is its ability to efficiently gather both local and global contextual information, resulting in heightened accuracy and diminished computational complexity. Notably, the Inception Network model relieves users of the burden of manually deciding on kernel selection or the necessity of pooling layers. The network autonomously determines these parameters. Users can input a comprehensive set of potential values for these parameters into the network, allowing it to autonomously learn the optimal parameters and convolution kernel combinations. Consequently, inception modules can be stacked iteratively to construct a more extensive network, effectively expanding both its depth and width.

### v2

@@ Abstract
Keywords: AI-assisted mathematics, Machine learning, Calabi-Yau manifolds, String compactification, Statistics and geometry
@@ 1 Introduction
In the realm of contemporary scientific inquiry, the application of deep learning has emerged as a transformative force, endowing researchers with novel tools and methodologies to explore intricate scientific phenomena. While AI has been transformative in the experimental sciences for decades, its revolution in theoretical and mathematical sciences have been a relatively more recent emergence (He:2017aed; Krefl:2017yox; Carifio:2017bov; Ruehle:2017mzq); see reviews and prospects in He:2018jtw; Ruehle:2020jrk; He:2021oav; Bao:2022rup; he2024can; gukov2024rigor; He:2024gnk. A programme to machine-learn various structures in mathematics by looking at pure mathematical data has been launched over the past seven years, and has blossomed to include algebraic geometry (He:2017aed; He:2018jtw), algebra and representation theory (He:2019nzx; davies2021advancing), graph theory and combinatorics (He:2020fdg), knot theory (Craven:2020bdz; Gukov:2020qaj; davies2021advancing), number theory (Alessandretti:2019jbs; He:2022pqn), symbolic computation (peifer2020learning; England_2018; lample2019deep), etc.
@@
Calabi-Yau manifolds are a protagonist in modern mathematics, residing at the interface between algebraic and differential geometry, mathematical physics (especially string theory), topology, number theory, and dynamical systems. These are Kähler manifolds that admit zero Ricci curvature. The simplest example is one that is familiar to the beginning student: the torus $T^{2}=S^{1}\times S^{1}$ can be complexified into a Riemann surface, which is in turn trivially Kähler and furthermore is Ricci-flat. The name “Calabi-Yau” originated in the Fields-winning proof by one of the authors (STY) yau1977calabi that settled the Calabi Conjecture (calabi1957kahler), and was first coined by physicists candelas1985vacuum.
@@
In searching for the Standard Model of particle physics within string theory compactifications, one of the first datasets in modern geometry was created; this was the so-called CICYs, for “Complete Intersection Calabi-Yau threefolds” (candelas1988complete; gagnon1994exhaustive; green1989all) (for pedagogical introductions, see the classic in hubsch1992calabi, as well as a recent treatment He:2018jtw). These are Calabi-Yau manifolds of complex dimension three, realized as algebraic varieties embedded as complete intersections in products of projective spaces. From a data-scientific point of view, they are rather straight-forwardly represented as matrices of non-negative integers, which encode the multi-degree information of the homogeneous polynomials that define the variety. It was natural, therefore, that the first machine-learning experiments in pure mathematics was in consideration on this data-set (He:2017aed). Indeed, despite its rich properties, the precise nature of the Calabi-Yau manifold often proves challenging to intuitively grasp and compute, presenting an intriguing and demanding avenue for the application of deep learning.
@@
Now, one should bear in mind that in topology, characteristics that describe the structure of manifolds, such as the classic Betti numbers which describe the closed cycles, are often intricately and closely linked with the data representing these manifolds. This makes the precise derivation of these characteristics quite challenging. The analytic mathematical formulas for these properties are largely unexplored  [^1] Due to the moduli dependence of quantities such as cohomology groups—unlike their alternating sums, which give rise to Euler characteristics that can be more readily computed by the index theorem—they are very difficult to calculate. See Constantin:2018hvl for a recent successful extraction of exact formulae for surfaces guided by machine learning., and their computation relies on algorithms that are both complex and time-consuming. Therefore, the application of deep learning in the field of topology poses significant challenges. This paper is dedicated to utilizing the capabilities of deep neural networks to predict some key properties of the Calabi-Yau manifold, in order to explore the potential applications of deep learning within the realm of mathematics and theoretical science.
@@
Henceforth, we focus on CICY threefolds, which have been a pivotal class of manifolds in the context of string model building and, as mentioned, one of the first data-set in pure mathematics to enter the age of data (candelas1988complete) and the era of machine-learning (He:2017aed). Since then, there has been a host of activity on machine-learning properties of CICYs. Neural network methodologies were applied to predict the Hodge numbers of CICY three-folds, favorability, and discrete symmetries (bull2018machine; Bull:2019cij), indicating that they outperform classical statistical methods such as Support Vector Machines (SVM). Subsequently, the prediction accuracy of Hodge numbers were increased to over 99% by leveraging the Inception network model (erbin2021inception; Erbin:2020tks), and more recent comparative study, with impressive accuracies, was performed in Keita:2024skh. Other related directions have included CICY fourfolds (he2021machine; Erbin:2021hmx) and generalized CICYs (Cui:2022cxe) (see review in Erbin:2023ncy), numerical metrics (Ashmore:2019wzb; Douglas:2020hpv; Anderson:2020hux; Anderson:2023viv; Ashmore:2023ajy), weighted CICYs (Berman:2021mcw; Hirst:2023kdl; macfadden2024dna), distinguishing elliptic fibrations (He:2019vsj) and topologies (Jejjala:2022lxh; chandra2023manifolds; gendler2023counting), etc. This expansion reflects a continued effort to broaden the scope of neural network applications in predicting geometric properties, highlighting the versatility of these models in handling increasingly complex mathematical structures.
@@
Now, most of the work on machine-learning of CICY topological invariants have focused on Hodge numbers, whereas, as we will soon see, more refined invariants involve Chern classes and intersection numbers. This is an important issue because the “topological type” of a (smooth, compact, and simply connected) Calabi-Yau threefold, due to the extension of a theorem of Wall wall1966classification, consists of a pair of Hodge numbers $(h^{1,1},h^{2,1})$, as well as (when expressed in an appropriate basis of the Kähler classes) the second Chern class $(c_{2})_{r}$ and the triple-intersection numbers $d_{rst}$ of the curve classes. It is a standing conjecture of one of the authors (STY), that in any complex dimension $n$, the possible topological types of a Calabi-Yau $n$-fold is finite.
@@
In sum, we hope to explore and expand the potential of deep learning to predict more critical properties of the CICY three-folds. In this study, we seek to predict the triple intersection numbers of CICY three-folds, combining ideas from He:2017aed; erbin2021inception; Jejjala:2022lxh, by an inception network V3 model. By amalgamating the nonlinear feature extraction capabilities of deep learning with the geometric properties of manifolds, we aim to enhance the deep learning model’s understanding of the CICY three-folds.
Summary
The organization of the paper is as follows. We begin in Section 2 with a review of the construction of the complete intersection Calabi-Yau (CICY) database, emphasizing on their topological invariants, especially the triple-intersection numbers. Then, in Section 3, we compare four models - SVM, random forest, XGBoost, and Google Inception, in predicting the four key quantities $(d_{1},d_{2},d_{3},d_{p})$ defined in (11) which capture CICY topology. In Section 4, we report the prediction accuracies of these quantities, with the main conclusion being that the Inception Network can reach around 90% accuracy in the standard 80-20 data cross-split. Finally, we conclude with outlook in Section 5.
@@ Acknowledgments
YHH is grateful to the STFC, UK for grant ST/J00037X/2 and a Leverhulme project grant. ZGY has been supported by Singapore Ministry of Education Tier 2 grant A-8001562-00-00 and Tier 1 grant A-8002931-00-00 at the National University of Singapore. ZGY also thanks the London Mathematical Society and the Royal Statistical Society for their partial support of his research trip to the UK.
@@ 2.1 Dataset
CICY manifolds are realized as a complete intersection of polynomials in a product of complex projective spaces (candelas1988complete; gagnon1994exhaustive). What this means is the following. Consider the ambient space as the product $X=\mathbb{C}P^{n_{1}}\times\cdots\times\mathbb{C}P^{n_{m}}$, a CICY three-fold $M^{3}$ is embedded as $k$ polynomials, the complete intersection condition is that
@@
Thus, each deformation family of CICYs is represented by an $m\times k$ matrix of non-negative (and indeed largely sparse) integers.
Now, $M^{3}$ is a Calabi-Yau three-fold, so it has vanishing first Chern class. This conveniently translates to (using adjunction, see e.g., hubsch1992calabi; He:2018jtw) the condition that row sums to one less than dimension of the corresponding projective factor:
@@
The CICY three-folds were classified in candelas1988complete, using the then state-of-the-art computing. Up to trivial permutation equivalence and also simple equivalence due to birational transformations, there are 7890 configuration matrices, giving rise to perhaps the first “big” database in geometry. Today, this database was updated in Anderson:2007nc and maintained at
@@
descends from the ambient $X$, and and we have, correspondingly, $k$ cohomology classes $\sum\limits_{i=1}^{m}q_{i}^{j}x_{i}$ for $j=1,\ldots,k$. It should be emphasized that this implies that while $h^{1,1}(M^{3})$ of these classes come at least from the $m$ projective factors, by no means they are equal in general, as extra Kähler classes arise from the restriction. In the case of equality $h^{1,1}(M^{3})=m$, when all divisors of $M^{3}$ descend from the simple ambient space $X$, we refer to $M^{3}$ as favourable. Of the original 7890 CICY threefold configuration matrices (candelas1988complete), there are 4896 favorable geometries and 2994 unfavorable geometries (anderson2017fibrations). Furthermore, new and favourable descriptions of 2946 unfavorable CICYs were presented therein. The favorable list can be obtained from http://www1.phys.vt.edu/cicydata/.
Here, we briefly summarize $(d_{1},d_{2},d_{3},d_{p})$ for all 7,680 samples with the empirical distributions:
  • $d_{1}$: takes values in $\{1,2,3,4,5,6,8,9,12,16\}$. The vast majority are $d_{1}=1$ (7336 samples), with a smaller cluster at $d_{1}=2$ (413). Other values are rare: e.g. $d_{1}=3$ (24), $d_{1}=4$ (34), and only isolated cases at $d_{1}=5,6,8,9,12,16$.
  • $d_{2}$: shares the same range $\{1,2,3,4,5,6,8,9,12,16\}$, but exhibits a different balance: $d_{2}=1$ (5315) and $d_{2}=2$ (2402) dominate, while higher values (3–16) occur with very low frequency.
  • $d_{3}$: has extended range $\{1,2,3,4,5,6,8,9,12,16,18\}$. Its distribution is sharply peaked at $d_{3}=2$ (3728) and $d_{3}=6$ (3710), almost evenly split. All other values appear in very small counts.
  • $d_{p}$: shows the broadest spread, covering $\{2,4,6,8,12,18,24,36,44,50,52,54,56,60,64\}$. It is, however, highly concentrated at $d_{p}=4$ (3724) and $d_{p}=12$ (3688). Smaller groups include $d_{p}=2$ (275), $d_{p}=6$ (25), $d_{p}=8$ (61), and a tail of rare larger values up to $64$.
@@
Theorem 2.1 (C. T. C. Wall (wall1966classification)).
@@
For Calabi-Yau threefolds, $c_{1}=0$ by definition. Furthermore, for the second Chern class, we can fix a Kähler basis $\{J^{r}\}_{r=1,2,\ldots,h^{1,1}}$ for $H^{2}(M^{3};\mathbb{Z})$, into which one can expand $c_{2}=\sum\limits_{s,t=1}^{h^{1,1}}(c_{2})_{st}J^{s}J^{t}$. Likewise, the triple intersection numbers record the information about $H^{1}(M^{3};\mathbb{Z})\times H^{1}(M^{3};\mathbb{Z})\times H^{1}(M^{3};\mathbb{Z})\longrightarrow\mathbb{C}$, can be expressed in this basis as the triple integral $d_{rst}=\int_{M^{3}}J^{r}\wedge J^{s}\wedge J^{t}$. Moreover, for simply-connected Calabi-Yau threefolds (with which we focus here, and to which class all CICY certainly belong), the only non-trivial Hodge numbers are $(h^{1,1},h^{2,1})$ (these were computed in green1989all). Therefore, the topological type of Calabi-Yau threefolds will be given by the following list of non-negative integers:
@@
This list of integer (tensors) is the topological datum for our Calabi-Yau threefold.
@@ 2.2 CICY Intersections
where the first equality follows from Gauss-Bonnet-Chern in expressing the Euler number in terms of the top Chern class, and the second equality come from expressing the Euler number in terms of an alternating sum over the Hodge (Betti) numbers. Also, by Poincaŕe duality, $H^{4}(M,\mathbb{Z})\cong H^{2}(M,\mathbb{Z})$, thus second Chern class $c_{2}$ can be labelled by $h^{1,1}$ as
$c_{2}(M)]_{t}=\sum_{r,s}[c_{2}(M)]_{rs}d_{rst}.$ (10)
@@
In general the computation of $(d_{rst},c_{2})$ is applicable for favorable geometries, we will design the computer algorithm later. On the other hand, it is usually hard to determine whether two sets of data $(d_{rst},c_{2})$ corresponding to the same topology, i.e., to determine whether the difference between them corresponding merely to a change of basis of $H^{2}(M,\mathbb{Z})$. Thus, we adopt a family of divisibilitiy invariants (see §8.1 of hubsch1992calabi) up to change of basis:
@@
will be used to distinguish CICY topology. Since Hodge number prediction had been the subject of the literature since 2017, we will here focus on $(d_{1},d_{2},d_{3},d_{p})$. To our knowledge, the compilation of these four quantities for manifolds is a new addition to existing databases.
@@ 2.3 Computer Algorithm for $d_{rst}$
    where $\sigma=(\sigma_{1},\sigma_{2},\cdots,\sigma_{m})$ is the partition of set $k]=\{1,\cdots,k\}$, $q_{i}^{\sigma_{i}}:=\prod_{j\in\sigma_{i}}q_{i}^{j}$. $|\sigma_{i}|=n_{i},\ i\neq r,s,t$; $|\sigma_{i}|=n_{i}-1$ otherwise.
@@
    Second, we construct a $k\times k$ square matrix $A_{rst}(i,j)$ by adding $n_{i}-1$ copies of the $i$-th row to $A$ if $i\neq r,s,t$. Otherwise, we add $n_{i}-2$ copies of the $i$-th row to $A$. Note that the subscript rstrst is just to show that the matrix depends on the indices of the triple intersection numbers $d_{rst}$ and the row-column indices that actually define the matrix are denoted as $(i,j)$.
@@
1:function detm(AA)
2:  if number of rows in A =1=1 then
@@
4:  else if number of rows in A =2=2 then
@@ 3.1 Inception Network model
To construct a robust architecture, we leverage the Inception Neural Network Model from GoogleNet (szegedy2016rethinking). This deep convolutional neural network architecture employs inception modules, a unique design that captures features across various spatial scales and complexities concurrently, departing from the sequential approach of traditional Convolutional Neural Networks. The distinctive feature of Inception is its ability to efficiently gather both local and global contextual information, resulting in heightened accuracy and diminished computational complexity. Notably, the Inception Network model relieves users of the burden of manually deciding on kernel selection or the necessity of pooling layers. The network autonomously determines these parameters. Users can input a comprehensive set of potential values for these parameters into the network, allowing it to autonomously learn the optimal parameters and convolution kernel combinations. Consequently, inception modules can be stacked iteratively to construct a more extensive network, effectively expanding both its depth and width.